OpenStudy (anonymous):

what are the factors of w^(4)-256?

jhonyy9 (jhonyy9):

(w^2 -16)(w^2 +16)=(w-4)(w+4)(w^2 +16)

OpenStudy (anonymous):

shouldn't the -16 in the first parenthesis be a +16?

jhonyy9 (jhonyy9):

no if you begin this calcules from formula a^2 -b^2 =(a-b)(a+b) - so than please you check it how i have wrote there are right sure